A stone is thrown upward with a velocity of 6.3m/s² from a bridge. If it falls after 2 seconds. Find the bridge height

Grade:11

1 Answers

Arun
6 years ago
u is given = 6.3 , total time of flight = 2sec
applying ,  S = ut + at2/2 ,
S = H = ? , a=-g , u=+6.3m/s (upward)
putting these in above eq
H = 6.3 *2 - 9.8*4/2 = 12.6 - 19.6 = 7 m
Regards
Arun (askIITians forum expert)